Problem 7 (5 points each) a./ (Inx) dx b. fx.ex+3dx c. ſcosx. 5sinx+3dx
Answer #1

(9) Let I = f (ln() e da but u= ln(a) du = – da Is a furdu ID# [1964] * +c] 1 fx. enos die X3 103. - put u=-x²+3 du = - za do